{2x + y=7
{-5 - 2y=8

Answer by tutorvolunteer(9)

This is a system of equations. There are many methods for solving this, i.e. substitution, plugging in.
I will solve this by plugging in Y to get X.
2x + y = 7, so y = 7-2x
Also, since -5-2y = 8, we can plug in y=7-2x, and get:
-5-2(7-2x) = 8
We distribute -2, so we get: -5 -14 + 4x = 8.
Simplifying this, we get: x=6.75
Plug this into y = 7 - 2x, and we get: y = - 6.5